### v4

# Restivo Salemi property for $\alpha$-power free languages with $\alpha\geq 5$ and $k\geq 3$ letters

Abstract
In 2009, Shur published the following conjecture: Let $L$ be a power-free language and let $e(L)\subseteq L$ be the set of words of $L$ that can be extended to a bi-infinite word respecting the given power-freeness. If $u,v\in e(L)$ then $uwv\in e(L)$ for some word $w$.
Let $L_{k,\alpha}$ denote an $\alpha$-power free language over an alphabet with $k$ letters, where $\alpha$ is a positive rational number and $k$ is positive integer. We prove the conjecture for the languages $L_{k,\alpha}$, where $\alpha\geq 5$ and $k\geq 3$.
Keywords: Transition Property; Power Free Words; Restivo Salemi Property
1 Introduction
Due to a significant overlap of notation, we copy and use the definitions and some explications from sections “Introduction” and “Preliminaries” from [6].
Let $r$ be a non-empty word, let $\alpha$ be a positive rational number, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$. We call the word $r^{\alpha}$ an $\alpha$-power of $r$. For example $(1234)^{3}=123412341234$, $(1234)^{\frac{7}{4}}=1234123$, and $(1234)^{\frac{1}{2}}=12$.
Suppose a finite or infinite word $w$. Let
$\begin{split}\Theta(w)=\{(r,\alpha)\mid r^{\alpha}\mbox{ is a factor of }w\mbox{ and }r\mbox{ is a non-empty word and }\\ \alpha\mbox{ is a positive rational number}\}\mbox{.}\end{split}$
We say that $w$ is $\alpha$-power-free if
$\{(r,\beta)\in\Theta(w)\mid\beta\geq\alpha\}=\emptyset$
and we say that $w$ is $\alpha^{+}$-power-free if
$\{(r,\beta)\in\Theta(w)\mid\beta>\alpha\}=\emptyset\mbox{.}$
The square free ($2$-power-free) and cube free words ($3$-power-free) are well known examples of the power-free words. More detailed introduction into the field of power-free words can be found, for instance, in [3] and [9].
In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power-free words [4]. Problem $4$ and Problem $5$ are relevant for the current article, hence we recall these two problems:
  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power-free.
  • Problem $5$: Given finite $\alpha$-power-free words $u$ and $v$, find a transition word $w$, if it exists.
In 2019, a survey on the solution of all the five problems has been presented in [2]; in particular, Problem $4$ and Problem $5$ were resolved for some binary languages. In addition, in [2] the authors resolved Problem $4$ and Problem $5$ for cube free words.
Let $\mathbb{N}$ denote the set of positive integers and let $\mathbb{Q}$ denote the set of rational numbers.
Definition 1.1.
([5, Definition $1$]) Let
$\begin{split}\Upsilon=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k=3\mbox{ and }\alpha>2\}\\ \cup\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k>3\mbox{ and }\alpha\geq 2\}\\ \cup\{(k,\alpha^{+})\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 2\}\mbox{.}\end{split}$
Remark 1.2.
([5, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a number with $+$ then $k\geq 3$ and $\alpha\geq 2$. If $(k,\alpha)\in\Upsilon$ and $\alpha$ is just a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
In 2020, Problem $4$ and Problem $5$ were resolved for $\alpha$-power-free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [5].
In 2009, a conjecture related to Problems $4$ and Problem $5$ of Restivo and Salemi appeared in [8]:
Conjecture 1.3.
Let $L$ be a power-free language and let $e(L)\subseteq L$ be the set of words of $L$ that can be extended to a bi-infinite word respecting the given power-freeness. If $u,v\in e(L)$ then $uwv\in e(L)$ for some word $w$.
In 2018, Conjecture 1.3 appeared also in [7] using a “Restivo Salemi property”; it was defined that a language $L$ has the Restivo Salemi property if Conjecture 1.3 holds for the language $L$. (Thus Restivo-Salemi property is exactly the transition property for extendable power-free languages.)
[…]
In [7, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$). Recall that $2^{+}$-power-free language is the overlap free language. Otherwise, as to our best knowledge, Conjecture 1.3 remains open.
Let
$\widetilde{\Upsilon}=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 5\}\mbox{.}$
Clearly $\widetilde{\Upsilon}\subset\Upsilon$. In the current article, we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$; see Corollary 3.10. The requirement that $\alpha\geq 5$ was necessary in the proof of Proposition $3$ in [6]. Since we use the results of [6], we require also for the current paper that $\alpha\geq 5$.
In order to prove our result, we apply some ideas and results from [5] and [6]. We briefly explain the very basic idea. If $u,v$ are square power-free words ($2$-power-free) and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free. Just note that there cannot be a factor in $uxv$ which is a square and contains $x$, because $x$ has only one occurrence in $uxv$.
This very basic idea was generalized to a finite number of occurrences of $x$ instead of a one single occurrence of $x$: In [5], the author constructed right and left infinite words containing a given factor and a non-recurrent letter $x$. The non-recurrence of the letter $x$ and several involved observations of recurrent and non-recurrent factors allowed the author to resolve Problem $4$ and Problem $5$ of Restivo and Salemi for languages $L_{k,\alpha}$, where $(k,\alpha)\in\Upsilon$.
In 2023, the construction of right and left infinite power-free words with a non-recurrent letter has been extended to bi-infinite $\alpha$-power-free words over an alphabet with $k$-letters, where $(k,\alpha)\in\widetilde{\Upsilon}$ [6]. In [6], it was also predicted that these bi-infinite power-free words with a non-recurrent letter could help to prove Conjecture 1.3. In the current article we apply indeed these bi-infinite $\alpha$-power-free words with a non-recurrent letter $x$ to prove Conjecture 1.3 for languages $L_{k,\alpha}$, where $(k,\alpha)\in\widetilde{\Upsilon}$.
2 Preliminaries
Let $\Sigma_{k}^{+}$ denote the set of all non-empty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
Let $\Sigma_{k}^{\infty}=\Sigma_{k}^{\mathbb{N},L}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{Z}}$. We call $w\in\Sigma_{k}^{\infty}$ an infinite word.
Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the non-empty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. If $w\in\Sigma_{k}^{\infty}$ and $\operatorname{occur}(w,t)=\infty$, then we call $t$ a recurrent factor in $w$.
Let $F(w)$ denote the set of all finite factors of a finite or infinite word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. The set $F(w)$ contains the empty word and if $w$ is finite then also $w\in F(w)$. Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent non-empty factors of $w\in\Sigma_{k}^{\infty}$.
Let $\operatorname{Prf}(w)\subseteq F(w)$ denote the set of all prefixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}$ and let $\operatorname{Suf}(w)\subseteq F(w)$ denote the set of all suffixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},L}$. We define that $\epsilon\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$ and if $w$ is finite then also $w\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$.
Let $L_{k,\alpha}^{\infty}\subseteq\Sigma_{k}^{\infty}$ denote the set of all infinite $\alpha$-power-free words over $\Sigma_{k}$; formally $L_{k,\alpha}^{\infty}=\{w\in\Sigma_{k}^{\infty}\mid F(w)\subseteq L_{k,\alpha}\}$. In addition we define $L_{k,\alpha}^{\mathbb{N},R}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},R}$, $L_{k,\alpha}^{\mathbb{N},L}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},L}$, and $L_{k,\alpha}^{\mathbb{Z}}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{Z}}$; it means the sets of right infinite, left infinite, and bi-infinite $\alpha$-power-free words, respectively.
We define the reversal $w^{R}$ of a finite or infinite word $w=\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{N},L}$ as follows:
  • $\epsilon^{R}=\epsilon$.
  • If $w\in\Sigma_{k}^{+}$ and $w=w_{1}w_{2}\dots w_{m}$, where $w_{i}\in\Sigma_{k}$ and $1\leq i\leq m$, then $w^{R}=w_{m}w_{m-1}\dots w_{2}w_{1}$.
  • If $w\in\Sigma_{k}^{\mathbb{N},L}$ and $w=\dots w_{2}w_{1}$, where $w_{i}\in\Sigma_{k}$ and $i\in\mathbb{N}$, then $w^{R}=w_{1}w_{2}\dots\in\Sigma_{k}^{\mathbb{N},R}$.
  • If $w\in\Sigma_{k}^{\mathbb{N},R}$ and $w=w_{1}w_{2}\dots$, where $w_{i}\in\Sigma_{k}$ and $i\in\mathbb{N}$, then $w^{R}=\dots w_{2}w_{1}\in\Sigma_{k}^{\mathbb{N},L}$.
Remark 2.1.
It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power-free word is also $\alpha$-power-free.
The next proposition is a reformulation of Corollary $1$ from [5] using only the notation of the current article.
Proposition 2.2.
(reformulation of [5, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
Remark 2.3.
([6, Remark $3$]) Proposition 2.2 says that if $z$ is a finite $\alpha$-power-free word that can be extended to a left infinite $\alpha$-power-free word having a letter $x$ as a recurrent factor and $s$ is a left infinite $\alpha$-power-free word not containing the letter $x$ as a factor, then there is a left infinite $\alpha$-power-free word containing $z$ as a suffix and having only a finite number of occurrences of $x$.
The next elementary lemma follows easily from famous theorems of Thue [10, 11].
Lemma 2.4.
(reformulation of [5, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let $\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\cup\{\epsilon\}\}\mbox{.}$ It means that all non-empty factors of $s\in\Phi(t)$ are recurrent factors of $t$. The following lemma is an immediate consequence of König’s lemma [1].
Lemma 2.5.
([5, Lemma $4$]) If $k\in\mathbb{N}$ and $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
The main result of [6] can be formulated as follows:
Theorem 2.6.
If $(k,\alpha)\in\widetilde{\Upsilon}$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma_{k}$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
Remark 2.7.
We will not apply Theorem 2.6 directly, because we will need some additional properties of the constructed bi-infinite word. Instead, we will apply [6, Theorem $1$] (see Theorem 3.4 in the current article).
3 Bi-infinite $\alpha$-power words
For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$. We define two technical sets $\Gamma$ and $\Delta$.
Definition 3.1.
([6, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
  • $w\in\Sigma_{k}^{+}$, $\eta,u\in\Sigma_{k}^{*}$, and
  • if $|u|\leq|w|$ then $|\eta|\geq(\alpha+1)\alpha^{|w|-|u|}|w|\mbox{.}$
Remark 3.2.
([6, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is non-empty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
Definition 3.3.
([6, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that
$(s,\sigma,w,\eta,x,u)\in\Delta$
if and only if
  1. 1.
    $s\in\Sigma_{k}^{\mathbb{N},L}$, $\sigma,\eta,u\in\Sigma_{k}^{*}$, $w\in\Sigma_{k}^{+}$, $x\in\Sigma_{k}$,
  2. 2.
    $s\sigma w\eta xu\in L_{k,\alpha}^{\mathbb{N},L}$,
  3. 3.
    $(w,\eta,u)\in\Gamma$,
  4. 4.
    $\operatorname{occur}(s\sigma w,w)=1$, and
  5. 5.
    $x\not\in F(s)\cup F(u)$.
In [6], it was shown that if $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$ and $t$ is a right infinite $\alpha$-power-free word with no occurrence of the letter $x$ then there is a bi-infinite $\alpha$-power-free word containing the factor $w$ and having only a finite number of occurrences of $x$:
Theorem 3.4.
([6, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
Definition 3.5.
([6, Definition $4$]) Suppose $v\in L_{k,\alpha}^{\mathbb{Z}}$ and $w\in F(v)\setminus\{\epsilon\}$. If there are $v_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $v_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ such that $v=v_{1}v_{2}$ and $w\in F_{r}(v_{2})$ then we say that $w$ is on-right-side recurrent in $v$.
Remark 3.6.
[6, Remark $8$]) Note in Definition 3.5 that no restriction is imposed on the recurrence of $w$ in $v_{1}$. This means that $w$ may also be recurrent in $v_{1}$.
We show how to construct an element of $\Delta$ from a bi-infinite $\alpha$-power-free word $v$ with a finite factor $w$ containing the letter $x$.
Lemma 3.7.
If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, and $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
Proof.
We distinguish following cases:
  1. 1.
    If $\operatorname{occur}(v,w)=\infty$ or $x$ is on-right-side recurrent in $v$ then it is easy to see that there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$, and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
  2. 2.
    Otherwise; it means that $\operatorname{occur}(v,w)<\infty$ and $x$ is not on-right-side recurrent in $v$. Then there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$,
    $x\not\in F(\widehat{v})\mbox{, }$ (1)
    and
    $\operatorname{occur}(\overline{v}w,w)=1\mbox{.}$ (2)
    Let $\eta_{0}\in\operatorname{Prf}(\widehat{v})$ be such that
    $|\eta_{0}|>\max\{|w\overline{\eta}|,(\alpha+1)\alpha^{|w|}|w|\}\mbox{.}$ (3)
    We claim that $\overline{v}w\overline{\eta}\eta_{0}x\in L_{k,\alpha}^{\mathbb{N},L}$. To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}\eta_{0}x)$. From (1) we have that $x\not\in F(\eta_{0})$; consequently $\eta_{0}x\in L_{k,\alpha}$ and $\eta_{0}x\in\operatorname{Suf}(r)$. From (3) we have that $|\eta_{0}|>|w\overline{\eta}|$; it follows that $w\in F(rr)$. From $\beta>\alpha\geq 5$ and (2), we conclude that are not such $r$ and $\beta$. Set $\eta=\overline{\eta}\eta_{0}$. From (3) it is obvious $(w,\eta,\epsilon)\in\Gamma$.
In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$, and $(w,\eta,\epsilon)\in\Gamma$.
  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is obvious that $(\overline{s},\epsilon,w,\eta,x,\epsilon)\in\Delta$.
  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $(k,\alpha)\in\widetilde{\Upsilon}$, Lemma 2.4 asserts that $t$ exists. Let $s=t^{R}$. Clearly $s$ is a left infinite $\alpha$-power-free word and $x\not\in F(s)$.
    Let $z=w\eta x$. Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z\in\operatorname{Suf}(su)$. Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma,\eta_{0}\in\Sigma_{k}^{*}$ such that $\overline{s}\sigma w\eta_{0}x=su$, $\eta\in\operatorname{Suf}(\eta_{0})$, $w\eta x\in\operatorname{Suf}(\overline{s}\sigma w\eta_{0}x)$, and $(\overline{s},\sigma,w,\eta_{0},x,\epsilon)\in\Delta$. Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$ and $\eta_{0}$.
This completes the proof.
Suppose two bi-infinite $\alpha$-power-free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $\widetilde{w}_{1},\widetilde{w}_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $\widetilde{w}_{2}$. We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power-free word.
Proposition 3.8.
If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Proof.
Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix of $s_{2}$ such that
$|p|>\max\{|\widetilde{w}_{1}|,|\widetilde{w}_{2}|\}\mbox{.}$ (4)
Let $h\in\operatorname{Prf}(t_{1})$ be the shortest prefix of $t_{1}$ such that $hp\in\operatorname{Prf}(t_{1})$ and $|h|>|p|$; such $h$ exists, because $F(s_{2})\subseteq F_{r}(t_{1})$. We have that $s_{1}\widetilde{w}_{1}hp\in L_{k,\alpha}^{\mathbb{N},L}$, since $hp\in\operatorname{Prf}(t_{1})$ and $s_{1}\widetilde{w}_{1}t_{1}\in L_{k,\alpha}^{\mathbb{Z}}$. We show that $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$. Since $p\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
Let $\widehat{w}_{2}\in\Sigma_{k}^{*}$ be such that $\widehat{w}_{2}x\in\operatorname{Prf}(\widetilde{w}_{2})$ and $x\not\in F(\widehat{w}_{2})$. Clearly $\widehat{w}_{2}$ exists and is uniquely determined; realize that $\widehat{w}_{2}$ may be the empty word.
Since $p\widehat{w}_{2}x\in F(r^{\beta})$ and $x\not\in F(p\widehat{w}_{2})$, we have that
$p\widehat{w}_{2}x\in F(rr)\mbox{.}$ (5)
Recall $x\not\in F(s_{1})\cup F(hp)\cup F(t_{2})$. Then it follows from (4) that
$\operatorname{occur}(s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (6)
From (5), (6), and $\beta>\alpha\geq 5$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
This completes the proof. ∎
The main theorem of the article says that if $w_{3},w_{4}$ are finite non-empty factors of bi-infinite $\alpha$-power-free words then there is a bi-infinite $\alpha$-power-free word $v$ and a finite word $w_{0}$ such that $w_{3}w_{0}w_{4}$ is a factor of $v$.
Theorem 3.9.
If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{3}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{4}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w_{0}\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{3}w_{0}w_{4}\in F(v)$.
Proof.
If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{3}\in F(v_{11})$, and $w_{4}\in F(v_{22})$. It is clear $v_{11}{v_{22}}$ is $\alpha$-power-free since $F(v_{11})\cap F(v_{22})=\emptyset$. The theorem follows.
Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$. Let $w_{1}\in F(v_{1})$ and $w_{2}\in F(v_{2})$ be such that $x\in F(w_{1})\cap F(w_{2})$, $w_{3}\in F(w_{1})$, and $w_{4}\in F(w_{2})$. It is clear that such $w_{1},w_{2}$ exist.
Lemma 3.7 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x,\epsilon)\in\Delta$. Lemma 3.7 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x,\epsilon)\in\Delta$.
Let $t_{1}\in L_{k-1,\alpha}^{\mathbb{N},R}$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $\alpha\geq 5$, Lemma 2.4 asserts that $t_{1}$ exists. Let $s_{2}\in\Phi(t_{1})$; Lemma 2.5 implies that $\Phi(t_{1})\not=\emptyset$ and consequently $s_{2}$ exists. Clearly $s_{2}$ is $\alpha$-power-free and $x\not\in F(s_{2})$.
Theorem 3.4 implies that there are $\widehat{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\widehat{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\widehat{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\widehat{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$. Since the reverse operation preserves the power-freeness we have also that $s_{2}x\widehat{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\widehat{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\widehat{\eta}_{2}w_{2}\sigma_{2}$. Then Proposition 3.8 implies that there is $w\in\Sigma_{k}^{*}$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Table 1 shows the structure of the resulting word using also the notation from the proof of Proposition 3.8.
$\widetilde{w}_{1}$ $\widetilde{w}_{2}$
$s_{1}$ $\sigma_{1}$ $w_{1}$ $\widehat{\eta}_{1}$ $x$ $h$ $p$ $x$ $\widehat{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
Table 1: The structure of the word $v$.
Since $w_{3}\in F(w_{1})$, $w_{4}\in F(w_{2})$, $w_{1}\in F(\widetilde{w}_{1})$, and $w_{2}\in F(\widetilde{w}_{2})$, the theorem follows. This completes the proof. ∎
Theorem 3.9 has the following obvious corollary.
Corollary 3.10.
Conjecture 1.3 holds for the languages $L_{k,\alpha}$, where $\alpha\geq 5$ and $k\geq 3$.